The average height of students in a freshman class of a certain school has been 168.67 cm with a population standard deviation of 8.67 cm. Is there a reason to believe that there has been a change in the average height if a random sample of 49 students in the present freshman class has an average height of 155.04 cm? Use a 0.1 level of significance.


What are the given? Write only the number. :

population mean:Blank 1 cm

population standard deviation:Blank 2 cm

sample size:Blank 3

sample mean:Blank 4 cm

level of significance:Blank 5


What are the critical values? Write the positive critical value first then the negative.

z:Blank 6 andBlank 7


What is the value of the calculated z? Round your answer to the nearest hundredths.

z:Blank 8

1
Expert's answer
2022-05-30T16:03:41-0400

population mean: "168.67" cm

population standard deviation: "8.67" cm

sample size: "49"

sample mean: "155.04" cm

level of significance: "0.1"


"z_c:" "-1.6449" and "1.6449"


The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{155.04-167.67}{8.67/\sqrt{49}}"

"\approx-10.20"

z: "-10.20"